# Comparing the information content of probabilistic representation spaces

### v2

@@ 3.2 Quick estimation based on fingerprinting a representation space
3.3 Discovering consistently learned information fragments
In the context of disentanglement, we are interested in the similarity of information contained in individual dimensions across an ensemble of models. We compute the pairwise similarities between all dimensions of all models, and then use density-based clustering to identify information content that is found repeatedly. OPTICS (Ankerst et al., 1999) traverses elements in a set according to proximity and produces a “reachability” profile that is similar to a dendrogram produced by hierarchical clustering but more comprehensible for large sets of points (Sander et al., 2003). Valleys indicate denser regions and natural groupings of elements in the set. The consistency of information fragments in individual latent dimensions can then be visualized with the OPTICS reachability profile, the reordered pairwise similarity matrix—which will show a block diagonal form if the fragments are highly consistent—and the NMI with ground truth generative factors, if available.
@@ 3.4 Model fusion
Consider a set of representation spaces found by an ensemble of weak learners. In the spirit of “knowledge reuse” (Strehl & Ghosh, 2002) originally applied to ensembles of hard clusters, we might obtain a superior representation space from the synthesis of the set. In contrast to other assessments of representation space similarity (Duan et al., 2020; Kim & Mnih, 2018; Duong et al., 2023), the proposed measures of similarity are composed of mutual information terms, which can be optimized with differentiable operations by a variety of means (Poole et al., 2019). We optimize a synthesis space by maximizing its average similarity with a set of reference spaces, performing gradient descent directly on the encoding of the data points used for the Bhattacharyya matrices. Instead of requiring many models to remain in memory during training of a new encoder, the Bhattacharyya matrices are computed for the ensemble once and then used for comparison during training, and training is quick because of the involved array operations.
Figure 3: Channel similarity of ensembles of fifty models. (a) The channel similarity found in ensembles of models on dsprites, for $\beta$-VAE and InfoGAN-CR. (b) For a $\beta$-VAE ensemble trained on cars3d, we retrieve the most central representation subspace to each cluster and perform latent traversals. Grouping v can be found in Appx. A. (c) We repeat the analysis for $\beta$-TCVAE and FactorVAE, and compare the information content of the reordered channels to the representatives from panel b. (d,e) Channel similarity and latent traversals for $\beta$-VAEs trained on fashion-mnist and celebA.
@@ 4.1 Comparison of related methods on synthetic spaces
We first evaluated the similarity of synthetic representation spaces using ours and related methods (Fig. 2). A dataset of 64 points were embedded to one- and two-dimensional representation spaces, and then the pairwise similarities between the spaces were computed. We evaluated NMI and VI with a Monte Carlo approach and the significantly faster Bhattacharyya fingerprint-based approach. We compare to a modification of CKA (Kornblith et al., 2019) that uses the Bhattacharyya matrices as the similarity measure between representations, and to stochastic shape metrics (Duong et al., 2023). For this small dataset, the Bhattacharyya estimates of NMI and VI are nearly the same as the Monte Carlo alternatives while offering a speed up of 100×100$\times$100 ×. CKA and NMI find comparable similarity structure between the spaces; while CKA lacks the information theoretic underpinnings of NMI, it offers an easy-to-use alternative by simply replacing a point-based distance with a statistical distance between representations. NMI and VI largely agree, though not always: representation space c is more similar to space f than space e according to NMI, and the reverse is true for VI. The magnitude of VI is tied to the magnitude of transmitted information, so when the information of the spaces drops, VI will uniformly assign high similarity. NMI is less dependent on the magnitude of communicated information, though it becomes more sensitive to the uncertainty of information estimates when magnitudes are small (discussed below).
@@
All methods except stochastic shape metrics detect the similarity of the quasi-one-dimensional encodings (representation spaces a, d, and f). These encodings are not related by simple transformations, though a downstream neural network would be able to extract similar information from each. Representation spaces g and i are similar by information content, as variations of splitting the data into two groups, but they are deemed different by stochastic shape metrics. While stochastic shape metrics satisfy the desirable properties of a metric, the generalized NMI/VI offer a more flexible sense of similarity, owing to their utilization of mutual information.
@@ 4.2 Unsupervised detection of structure: channel similarity
We next analyzed the consistency of information fragmentation in ensembles of generative models trained on image datasets. Every latent dimension (channel) of every model in an ensemble was compared pairwise to every other; we used the Bhattacharyya fingerprint approach with a sample size of 1000 images randomly selected from the training set. Before using OPTICS to group latent dimensions by similarity, we removed dimensions transmitting less than 0.01 bits of information. We found NMI to more successfully detect channels with similar information content, and use it for all results in this section (comparison with VI in Appx. D).
In Fig. 3, the matrices display the pairwise NMI between all informative dimensions in the ensemble, and have been reorganized by OPTICS such that highly similar latent dimensions appear as blocks along the diagonal. On the left of each similarity matrix is the NMI with the ground truth generative factors (or other label information), and above is the reachability profile where identified groupings of repeatedly learned channels are indicated with shading and Roman numerals.
@@ 4.3 Assessing the content of the full latent space
If an ensemble of models fragments information into channels inconsistently (e.g., Fig. 3a), is it because the information content of the full latent space varies across repeated runs? We compared the information contained in the full 10-dimensional latent spaces of five models from each VAE-variant ensemble (Fig. 4). The amount of information was generally too large for estimation via Bhattacharyya matrices, which saturates at the logarithm of the fingerprint size, so we used Monte Carlo estimates of the mutual information (Appx. E). Whereas NMI proved more useful for the channel similarity analysis, VI tended to be more revealing about the heterogeneity of full latent spaces. Due to its normalization, NMI becomes dominated by uncertainty when the information contained in the latent space $I(U;X)$ is small (error bars in Fig. 4 are the propagated uncertainty from the mutual information measurements).
Across several methods and several hyperparameters for each method (Fig. 4a; models were from Locatello et al. (2019)), the information contained in the full latent space was largely consistent over different initializations. For the cars3d dataset, the latent spaces nearly all conveyed the full entropy of the dataset, meaning all representations were well-separated for the dataset of almost 18,000 images; consequently, the information content was always highly similar. By contrast, none of the models trained on smallnorb contained more than around 80% of the information about the dataset, and the similarity of latent spaces across members of an ensemble was more dependent on method and hyperparameter. The consistency of repeat runs of $\beta$-VAE and $\beta$-TCVAE was nonmonotonic with $\beta$, whereas FactorVAEs varied more with increasing $\gamma$.
How does the consistency of latent space information evolve over the course of training? For $\beta$-VAE ensembles, we again compared the full latent spaces for five models with random initializations (Fig. 4b). For all datasets considered (cars3d, smallnorb, celebA), consistency across the models dropped at around the same point in training that total information increased dramatically; after this point, all models encapsulated nearly the same information in their latent spaces up to convergence. Interestingly, the magnitude of the drop in NMI (or increase in VI) was monotonic with $\beta$ for smallnorb while the nonmonotonic trend in consistency found in Fig. 4a was recovered by the end of training.
@@ 4.4 Model fusion in a toy example
Finally, we created a toy example to demonstrate the capacity for model fusion with the proposed measures of representation space similarity. Consider a dataset with a single generative factor with SO(2) symmetry, such as an object’s hue. Difficulties arise when the global structure of a degree of freedom is incompatible with the latent space (Falorsi et al., 2018; Zhou et al., 2019; Esmaeili et al., 2024). Here we focus on a trivial mismatch to clearly demonstrate the synthesis of information from multiple representation spaces.
@@
We trained an ensemble of $\beta$-VAEs, each with a one-dimensional latent space that was insufficient to represent the global structure of the generative factor. Fig. 5a shows an example latent space and its associated Bhattacharyya coefficient distinguishability matrix between posterior distributions. The matrix shows the two flaws in the latent space where similar values of the generative factor have dissimilar representations. Assuming the flaws are randomly distributed from one training run to the next—which need not be true—the fusion of several such latent spaces might yield an improved representation of the generative factor.
@@
We performed gradient descent directly on posterior distributions in a two-dimensional latent space, with the objective to maximize average similarity with the ensemble. Namely, we computed either NMI or the $\exp(-\text{VI})$ using the Bhattacharyya matrix corresponding to the optimizable representations—recalculated every training step—and those of the ensemble. As the ensemble size grew, the synthesized latent space more closely captured the global structure of the generative factor (Fig. 5b,c). We quantify the performance with the continuity metric used in Esmaeili et al. (2024), which was adapted from Falorsi et al. (2018), except that we used the Bhattacharyya distance between posteriors instead of the Euclidean distance between point-based representations (Appx. E).
@@
Both NMI and VI boosted an ensemble of weak representation spaces to represent the generative factor, and fidelity (as measured by continuity) increased for larger ensembles. By comparison, simply maximizing the average mutual information between the synthesis space and the ensemble (Fig. 5c, right), $\langle I(U;V_{i})\rangle_{i}$, scattered all representations such that $I(X;U)\approx H(X)$. The normalization terms of NMI and VI help to maintain the relational structure between data points. Finally, we note that after training the ensemble of weak models, neither the original data nor the models were needed to train the synthesis space: the latent representations were optimized directly from the Bhattacharyya fingerprints.
@@
Figure 5: Learning a factor with $SO(2)$ structure by fusing weak representation spaces. (a) A one-dimensional latent space of a $\beta$-VAE trained on a dataset generated from a single periodic factor that we visualize with color hue. Flaws can be seen in the arrangement of posterior distributions (left) and the distinguishability matrix of Bhattacharyya coefficient values between posteriors, $\text{BC}_{ij}$ (right). (b) We optimized a synthesis representation space to have information content that maximizes similarity with an ensemble of one-dimensional spaces such as the one in panel a. The continuity of statistical distances between neighboring points, a proxy for the fidelity of the global structure of the generative factor, is shown for VI and NMI with slight horizontal offsets for visibility. Error bars are the standard deviation over five repeats of the experiment. (c) Synthesized representation spaces, where posterior means are displayed as points and the covariances as shaded ellipses, and their corresponding distinguishability matrices, where the average NMI (left, middle) and mutual information (right) were maximized.
@@ 5 Discussion
The processing of information can be directly assessed when representation spaces are probabilistic because information theoretic quantities are well-defined. Here we built upon a perspective of probabilistic representation spaces as soft clustering and proposed a natural generalization of two classic measures to compare clustering assignments. Originally, NMI and VI used the entropy of hard clustering assignments to normalize mutual information, but because $H(U)=I(X;U)=I(U;U^{\prime})$ for a hard clustering $U$ of data $X$, any of the three quantities could have been used. Only the third quantity, $I(U;U^{\prime})$, maintains the desirable standardization of $\text{NMI}=1$ and $\text{VI}=0$ for identical soft clustering assignments. By focusing on the information content of a representation space, comparisons are agnostic to aspects of the spaces such as their dimensionality, their discrete or continuous nature, and even whether a space is side information attached to the dataset in the form of labels or annotations (Newman & Clauset, 2016; Savić, 2018; Bazinet et al., 2023).
A more subtle contribution of this work is to shift the current focus of unsupervised disentanglement evaluation. As is clear from the cars3d information fragmentation (Fig. 3b), existing metrics that compare latent dimensions to ground truth generative factors can completely miss consistent information fragmentation. Unsupervised methods of evaluation (Duan et al., 2020; Lin et al., 2020) assess consistency of fragmentation at the scale of models, which can obscure much about the manner of fragmentation. Consider Fig. 3e, where two fragments of information about the celebA dataset are significantly more robust than others—what is special about these two pieces of information? We argue that more fine-grained inspection of information fragmentation is essential for a deeper understanding of disentanglement in practice.
The current work largely focused on comparing the information content across repeat training runs in an ensemble, making computational costs an important consideration. We found 50 models per ensemble to be sufficient to assess channel similarity, and only 5 models per ensemble for the full latent space; further, we found that there is still much to learn from ensembles of relatively simple models. On our machine with public code, training a single model from a recently proposed method (QLAE, Hsu et al. (2023)) took more than five hours, whereas we could train ten $\beta$-VAEs with a simpler architecture in the same time. Model fusion with weak models that are inexpensive to train, as in Sec. 4.4, might offer a promising alternative to representation learning with more computationally expensive models.

### v3

@@ 3.2 Routes to estimation
3.3 Discovering consistently learned information fragments via OPTICS clustering of latent dimensions
In the context of disentanglement, we are interested in the similarity of information contained in individual dimensions across an ensemble of models. We compute the pairwise similarities between all dimensions of all models—for example, 10 dimensions each from 50 models in an ensemble for a 500$\times$500 matrix of similarity values in Sec. 4.2—and then use density-based clustering to identify information content that is found consistently. OPTICS (Ankerst et al., 1999) is an algorithm that orders elements in a set according to similarity and produces a “reachability” profile where natural groupings of elements are indicated through valleys. An OPTICS profile is analogous to a dendrogram produced by hierarchical clustering, but can be more comprehensible for large sets of points (Sander et al., 2003). The consistency of information fragments in the latent dimensions of an ensemble of representation spaces can then be visualized with the OPTICS reachability profile, the reordered pairwise similarity matrix—which will have block diagonal form if the fragments are consistent—and the information contained about ground truth generative factors, if available.
@@ 3.4 Model fusion
Consider a set of representation spaces found by an ensemble of weak learners. In the spirit of “knowledge reuse” (Strehl & Ghosh, 2002) originally applied to ensembles of hard clusters, we might obtain a superior representation space from the synthesis of the set. In contrast to other assessments of representation space similarity (Duan et al., 2020; Kim & Mnih, 2018; Duong et al., 2023), the proposed measures of similarity are composed of mutual information terms, which can be optimized with differentiable operations by a variety of means (Poole et al., 2019). We optimize a synthesis space by maximizing its average similarity with a set of reference spaces, performing gradient descent directly on the encoding of the data points used for the Bhattacharyya matrices. Instead of requiring many models to remain in memory during training of a new encoder, the Bhattacharyya matrices are computed for the ensemble once and then used for comparison during training, and training is fast because of the involved array operations.
Figure 2: Comparing similarity measures for synthetic embedding spaces. (a) A dataset of 64 points, $x=1,...,64$, is transformed into nine representation spaces marked i-ix. Each posterior distribution $p(u|x)$ is a Gaussian with diagonal covariance matrix and standard deviations indicated by the colored ellipses. (b) We trained a classification head on top of the latent spaces of a to predict the input $x$, given a sample from the posterior distribution $p(u|x)$. The predicted probability distributions $p(\hat{x}_{j}|x_{i})$ are displayed as a matrix, with row corresponding to the input $x_{i}$. (c) The pairwise distinguishability of data points $x_{i}$ and $x_{j}$, as computed by the Bhattacharyya coefficient, serves as a fingerprint of the information content of the latent space. (d) The pairwise similarity of the representation spaces in panel a, found by a variety of methods. Runtimes to calculate the full matrix are shown above each method, except for Jensen-Shannon divergence (JSD) because it required training an additional classification network on top of each latent space. The stochastic shape metric requires the dimensionality of the compared spaces to match; undefined entries are grayed out. The Spearman rank correlation between similarity measures is shown in the bottom right.
@@ 4.1 Comparison of related methods on synthetic spaces
We first evaluated the similarity of synthetic representation spaces using ours and related methods (Fig. 2). A dataset of 64 points was embedded to one- and two-dimensional representation spaces, each communicating different information about the dataset. To assess the information available for downstream processing, we trained a classification head for each of the nine latent spaces to predict the input $x$ from a sample of the corresponding posterior, $u\sim p(u|x)$. As shown in Fig. 2b, the classification head’s average predictions, $p(\hat{x}_{j}|x_{i})=\mathbb{E}_{u\sim p(u|x_{i})}[p(\hat{x}_{j}|u)]$, reflected the overlap of posterior distributions in latent space by assigning non-zero probabilities to multiple outputs. The structure of these output probabilities closely matched the similarity patterns in the distinguishability matrices of panel c, supporting our use of the latter as fingerprints of the information in the latent space. Unlike the classification head’s outputs, the distinguishability matrices were directly accessible from the latent space without additional network training.
We evaluated the pairwise similarity between the nine latent spaces of panel a using several methods (Fig. 2d). To assess the information content available for downstream processing, we compared the average predictions of the trained classification heads for each input $x_{i}$ through the Jensen-Shannon divergence, $\langle\text{JSD}(p^{\alpha}(\hat{x}|x_{i})||p^{\beta}(\hat{x}|x_{i})\rangle_{i}$. If classification heads for latent spaces $\alpha$ and $\beta$ produce similar output distributions, the latent spaces likely share similar information content. Next, CKA (Kornblith et al., 2019) can use different measures of representation similarity as its basis of comparison; we used the inner product between means of the posterior distributions for the representational similarity in linear CKA, an exponential function of the Euclidean distance for nonlinear CKA, and the Bhattacharyya coefficient between posteriors as a statistical basis of representational similarity. Nonlinear CKA requires selecting a distance parameter for each space: we used the arithmetic mean of the standard deviation of every representation. Next, we computed similarity according to stochastic shape metrics (Duong et al., 2023). Finally, we computed VI and NMI, both with a Monte Carlo approach and the significantly faster Bhattacharyya fingerprint-based approach.
Spearman’s rank correlation of the similarity values between all pairs of synthetic spaces reveals multiple relationships between our proposed measures and the baselines. First, for this small dataset, the Bhattacharyya estimates of NMI and VI are highly consistent with the Monte Carlo alternatives while offering a speed up of 100×100$\times$100 ×. Second, VI captures similarity much the same as when comparing the outputs of downstream classification heads (JSD) even though the former can be evaluated directly from the latent space. Third, NMI relates the latent spaces in much the same way as the Bhattacharyya variant of CKA. While this ad hoc variant of CKA lacks the information-theoretic underpinnings of NMI, it offers an easy-to-use alternative by simply replacing a point-based distance with a statistical distance between representations.
@@
We additionally observe differences between the methods in relative similarities between the latent spaces. All methods except linear CKA and stochastic shape metrics detect the similarity of the quasi-one-dimensional encodings (representation spaces i, iv, and vi). These encodings are not related by simple transformations, though a downstream neural network can extract similar information from each (see JSD). Representation spaces vii and ix are similar by information content, as variations of splitting the data into two groups, but they are deemed different by all but NMI, VI, and the Bhattacharyya-based CKA. While stochastic shape metrics satisfy the desirable properties of a metric, the generalized NMI/VI offer a measure of similarity more relevant for downstream processing owing to their utilization of mutual information.
Figure 3: Assessing the consistency of channel information in ensembles of models. We used NMI as a similarity measure for OPTICS to detect fragments of information that are consistently stored in individual channels in an ensemble of trained models. (a) The channel consistency of models trained on dsprites, for $\beta$-VAE and InfoGAN-CR. The information with respect to generative factors is shown on the left of each similarity matrix. The $\beta=4$ $\beta$-VAE fragmented information inconsistently compared to the other two ensembles. (b) For a $\beta$-VAE ensemble trained on cars3d, the information content of channels was highly consistent, with seven distinct combinations of the three generative factors. Latent traversals for a representative channel from each grouping visualize the information content. (c) We compare the information content of the representatives from panel b to that of channels in $\beta$-TCVAE and FactorVAE ensembles. (d,e) Channel similarity and latent traversals for $\beta$-VAE ensembles trained on fashion-mnist and celebA. Additional channel similarity analyses and latent traversals can be found in Appx. A.
@@ 4.2 Unsupervised detection of structure: channel similarity
We next analyzed the consistency of information fragmentation in ensembles of generative models trained on image datasets. Using fifty models with ten latent dimensions (channels) each, for a variety of datasets, methods, and hyperparameters (some released with Locatello et al. (2019) [^1] https://github.com/google-research/disentanglement_lib/tree/master), we assessed structure in an ensemble’s channels. Every latent dimension of every model in an ensemble was compared pairwise to every other; we used the Bhattacharyya fingerprint approach with a sample of 1000 images randomly selected from the training set. Before using OPTICS to group latent dimensions by similarity (described in Sec. 3.3), we removed dimensions transmitting less than 0.01 bits of information. We found NMI to more successfully detect channels with similar information content, and use it in this section (comparison with VI in Appx. E).
In Fig. 3, the matrices display the pairwise NMI between all informative dimensions in the ensemble, and have been reorganized by OPTICS such that highly similar latent dimensions are close together and appear as blocks along the diagonal. On the left of each similarity matrix is the NMI with the ground truth generative factors (or other label information), and above is the OPTICS reachability profile where identified groupings of consistently learned channels are indicated with shading and Roman numerals.
@@ 4.3 Assessing the content of the full latent space
If an ensemble of models fragments information into channels inconsistently (e.g., Fig. 3a), is it because the information content of the full latent space varies across runs? We compared the information contained in the full 10-dimensional latent spaces of five models from each VAE ensemble (Fig. 4). The amount of information was generally too large for estimation via Bhattacharyya matrices, which saturates at the logarithm of the fingerprint size, so we used Monte Carlo estimates of the mutual information. Error bars in Fig. 4 are the propagated uncertainty from the mutual information measurements, with additional details in Appx. C. Whereas NMI proved more useful for the channel similarity analysis, VI tended to be more revealing about the heterogeneity of full latent spaces. Due to the form of its normalization, NMI becomes unreliable—i.e., is plagued by large uncertainty—when the information contained in the latent space $I(U;X)$ is small.
Across methods and hyperparameters for each method, the information contained in the full latent space was largely consistent over an ensemble (Fig. 4a; models from Locatello et al. (2019)). For the cars3d dataset, most latent spaces conveyed the full entropy of the dataset, meaning all representations were well-separated for the dataset of almost 18,000 images, making the information content trivially similar. By contrast, none of the models trained on smallnorb contained more than around 80% of the information about the dataset, and the similarity of latent spaces across members of an ensemble was more dependent on method and hyperparameter. Ensembles of $\beta$-VAE and $\beta$-TCVAE were fairly consistent even though the transmitted information dropped considerably with increasing $\beta$; DIP-I (Kumar et al., 2018) showed the opposite behavior, where consistency varied strongly with $\lambda_{od}$ but transmitted information did not.
How does the consistency of latent space information evolve over the course of training? For $\beta$-VAE ensembles, we again compared the full latent spaces for five models with random initializations (Fig. 4b). For all datasets considered (cars3d, smallnorb, celebA), consistency across the models dropped at around the same point in training that total information increased dramatically; after this point, all models encapsulated nearly the same information in their latent spaces up to convergence.
@@ 4.4 Model fusion in a toy example
Finally, we demonstrate the capacity for model fusion with the proposed measures of representation space similarity. Consider a dataset with a single generative factor with SO(2) symmetry, such as an object’s hue. Difficulties can arise when the global structure of a degree of freedom is incompatible with the latent space (Falorsi et al., 2018; Zhou et al., 2019; Esmaeili et al., 2024). In this example, a one-dimensional latent space cannot represent the global circular topology of SO(2), leading to discontinuities in the representation.
@@
To clearly demonstrate the synthesis of information from multiple representation spaces, we trained an ensemble of $\beta$-VAEs, each with a one-dimensional latent space that was insufficient to represent the global structure of the generative factor. Fig. 5a shows an example latent space and its associated distinguishability matrix, showing pairwise Bhattacharyya coefficients between posterior distributions. The matrix reveals flaws in the latent space where similar values of the generative factor have dissimilar representations.
@@
Assuming the flaws are randomly distributed from one training run to the next—which need not be true—the fusion of multiple such latent spaces might yield an improved representation of the generative factor. We performed gradient descent directly on posterior distributions in a two-dimensional latent space, with the objective to maximize average similarity with the ensemble. Namely, we computed either NMI or the $\exp(-\text{VI})$ using the Bhattacharyya matrix corresponding to the optimizable representations—recalculated every training step—and those of the ensemble. As shown in Fig. 5b,c, the synthesized latent space more closely captured the global structure of the generative factor as the ensemble size grew. To quantify the performance, we employed the continuity metric used in Esmaeili et al. (2024) and adapted from Falorsi et al. (2018), substituting the Bhattacharyya distance between posteriors for the Euclidean distance typically used for point-based representations (Appx. F).
@@
Both NMI and VI proved effective in boosting an ensemble of weak representation spaces to represent the generative factor, with fidelity (measured by continuity) improving for larger ensembles. By comparison, directly maximizing the mutual information between the synthesis space and the ensemble, $\langle I(U;V_{i})\rangle_{i}$, resulted in scattered representations with no overlap (Fig. 5c, right). The normalization terms of NMI and VI were essential for preserving the relational structure between data points. Finally, it is worth noting that after training the ensemble of weak models, neither the original data nor the models were needed to train the synthesis space: the latent representations were optimized directly from the Bhattacharyya matrices.
@@
Figure 5: Fusing weak representation spaces. (a) Example of a one-dimensional latent space of a $\beta$-VAE trained on a dataset generated from a single periodic factor (color hue), which has SO(2) symmetry. The latent space exhibits flaws where similar values of the generative factor are mapped to dissimilar representations, as seen in the posterior distributions (left) and the distinguishability matrix of Bhattacharyya coefficients between posteriors, $\text{BC}_{ij}$ (right). (b) We optimized a synthesis representation space to maximize similarity with an ensemble of such one-dimensional latent spaces. The continuity of statistical distances between neighboring points, an assessment of the fidelity of the global structure of the generative factor, improved as the ensemble size grew. Error bars show the standard deviation over five experiments, and values are offset horizontally for visibility. (c) Synthesized two-dimensional representation spaces (posterior means shown as points; covariances as shaded ellipses) and their corresponding distinguishability matrices. Panels compare results when maximizing average NMI (left, middle) and mutual information (right).
@@ 5 Discussion
The processing of information can be directly assessed when representation spaces are probabilistic because information theoretic quantities are well-defined. In this work, we generalized two classic measures for comparing the information content of clustering assignments to make them applicable to probabilistic spaces. By focusing on the information content of a representation space, we can assess what information is available for downstream processing, while remaining agnostic to aspects of the spaces such as their dimensionality, their discrete or continuous nature, and even whether a space serves as auxiliary information like labels or annotations (Newman & Clauset, 2016; Savić, 2018; Bazinet et al., 2023). While the examples in this work focused on relatively low-dimensional latent spaces that are common in practice, scaling to higher-dimensional representation spaces may face challenges related to the reliable estimation of mutual information (McAllester & Stratos, 2020).
With differentiable formulations for NMI and VI, model fusion and the more general problem of representational alignment (Sucholutsky et al., 2023; Muttenthaler et al., 2024) can be effectively approached from an information theoretic perspective. Potential applications include aligning representation spaces across models trained on different subsets of data, improving ensemble methods, and evaluating consistency of representations in multitask learning or domain adaptation, where reconciling heterogeneous latent spaces is often crucial.
A more subtle contribution of this work is to shift the current focus of unsupervised disentanglement evaluation. As is clear from the cars3d information fragmentation (Fig. 3b), existing metrics that compare latent dimensions to ground truth generative factors can completely miss consistent information fragmentation. Unsupervised methods of evaluation (Duan et al., 2020; Lin et al., 2020) assess consistency of fragmentation at the scale of models, which can obscure much about the manner of fragmentation. Consider the fragmentation of information about the celebA dataset in Fig. 3e: an assessment of the similarity of models would find middling values across the ensemble and miss that two fragments of information are remarkably consistent. We argue that more fine-grained inspection of information fragmentation is essential for a deeper understanding of disentanglement in practice.
The current work largely focused on comparing repeat training runs in an ensemble, making computational costs a consideration. We found 50 models per ensemble to be sufficient to assess channel similarity, and 5 models per ensemble for the full latent space. We also found that there is much to learn from ensembles of relatively simple models. On our machine, training a single model from a recently proposed method (QLAE, Hsu et al. (2023)) took more than five hours, and in the same amount of time we could train ten $\beta$-VAEs with a simpler architecture. Model fusion with weak models that are inexpensive to train, as in Sec. 4.4, might offer a promising alternative to representation learning with more computationally expensive models.